# Capturing Minds, Not Just Words: Enhancing Role-Playing Language Models with Personality-Indicative Data

### v2

@@ Abstract
Role-playing agents (RPA) have been a popular application area for large language models (LLMs), attracting significant interest from both industry and academia. While existing RPAs well portray the characters’ knowledge and tones, they face challenges in capturing their minds, especially for small role-playing language models (RPLMs). In this paper, we propose to enhance RPLMs via personality-indicative data. Specifically, we leverage questions from psychological scales and distill advanced RPAs to generate dialogues that grasp the minds of characters. Experimental results validate that RPLMs trained with our dataset exhibit advanced role-playing capabilities for both general and personality-related evaluations. Code and data are available at this URL.
@@ 1 Introduction
While existing RPLMs well replicate knowledge and tones of the intended characters, they struggle with capturing their minds, in tasks such as personality assessment (Wang et al., 2024a) and decision simulation (Xu et al., 2024). This is partly because existing role-playing datasets focus on characters knowledge and tones (Wang et al., 2024b; Shao et al., 2023). However, capturing characters’ minds are crucial for developing authentic RPAs.
In this paper, we propose to develop RPLMs via personality-indicative data. Specifically, we collect these data through questions from psychological scales. These scale questions are designed to quickly capture broad aspects of personality traits in individuals. Hence, we leverage advanced RPAs for knowledge distillation from them. Then, we apply these data to develop RPLMs that better capture the minds of the intended characters.
Specifically, we construct a dataset RolePersonality based on questions from 14 different psychological scales, including both single-round and multi-round data, inspired by InCharacter (Wang et al., 2024a). The dataset encompasses a wide range of personality scales and dimensions, providing a comprehensive foundation for training RPLMs.
@@
Figure 1: The framework of building and utilizing RolePersonality. First, we obtain RolePersonality by distillation from advanced RPAs using scale questions. Then, we train RPLMs on RolePersonality to enhance their ability to capture characters’ minds.
@@ 2.1 Role-Playing Language Models
This method utilizes LLMs for generating conversations or human annotation to build and augment datasets. The topics come from corresponding literature (Shao et al., 2023), general task instructions (Wang et al., 2024b), and special scenarios such as personality tests (Wang et al., 2024a).
@@ 2.2 Construction of Role-Playing Agents
This approach fine-tunes a base model using custom or existing role-playing datasets. Shao et al. (2023); Yu et al. (2024) enhance foundation models with improved role-playing abilities using datasets featuring a variety of characters and scenarios. Zhou et al. (2023); Wang et al. (2024b) tailor LLMs to role-play specific characters.
@@
For more in-depth role-playing of a specific character, many efforts have focused on character-level engineering (Zhou et al., 2023; Wang et al., 2024b). They collect and process character-related data from corresponding sources, including collecting profile from Wikipedia (Shao et al., 2023). Typically, they add long-term memory to retrieve knowledge about the character based on similarity with user’s query (Li et al., 2023).
@@ 3.1 Dataset Construction
To simulate the deep thoughts underlying the characters, we generate persona-indicative data by utilizing psychological scale questions, inspired by InCharacter Wang et al. (2024a). In practice, we construct an RPA by pairing the RPLM with descriptions and the memory base of the target character (Li et al., 2023). RPAs are then engaged with open-ended questions derived from established psychological scales. These questions are designed to elicit the character’s mindset and behaviors in various scenarios. The questions were adapted from well-known psychological scales such as the Big Five Inventory (BFI) and 16Personalities. For more details, refer to Sec. A. We start by rewriting psychological scale questions and implementing a selection process to refine the data.
@@
Not all questions are suitable for all characters. A question that violates the character’s background may induce hallucinations. We introduce a filtering mechanism to exclude questions that do not fit the character’s background.
Scale Selection
Our scales are sourced from psychological scales (Bem, 1981; Barrick and Mount, 1991), utilizing the questions rewritten by InCharacter (Wang et al., 2024a). However, not all the scales are closely related to character personalities. We carefully selected a subset of these scales that best reflect character personality traits, forming the subset Part. The entire set of selected scales constitutes the subset Full.
Multi-Turn Dialogue
We incorporate multi-turn dialogues to maintain conversation consistency and enhance the model’s contextual understanding. We select questions from different dimensions within the same scale. These multi-turn data form subset Multi. The subset consisting of only single-round data is classified as subset Single.
3.2 Dataset Statistics
@@
4 Experiment
4.1 Settings
Fine-tuning
Baseline
To compare the effectiveness of different datasets, we fine-tune the model with the same settings on three subsets of RolePersonality (Full+Single, Part+Single, Part+Multi), the dataset introduced by CharacterLLM(Shao et al., 2023) and RoleBench(Wang et al., 2024b). For each dataset, we select approximately 20,000 samples for fine-tuning, keeping the data size about the same. These models, along with the original mistral-7B and gpt-3.5-turbo-0301, are subsequently evaluated to assess their performance.
Evaluation Protocols
@@ 4.2 Personality Fidelity (PF)
We use LLMs to judge the character’s personality based on the model’s responses to personality scale questions and compare the judgments with the ground truth, which was determined by human annotators. We select 8 test characters from the dataset proposed by InCharacter (Wang et al., 2024a). All data related to these test characters are excluded from the training set to ensure unbiased evaluation. This metric provides a comprehensive assessment of the model’s ability to accurately reflect a character’s holistic personality traits.
@@ 5 Conclusion
Our approach to developing Role-Playing Language Models (RPLMs) presents several risks. First, reliance on LLM-generated datasets may perpetuate inherent biases and inaccuracies, leading to unintended behaviors. Second, the lack of compliance mechanisms in interview data can result in inconsistencies, undermining authenticity. Third, the absence of human evaluation means subtle nuances in character portrayal may be missed by automated metrics. Ethical concerns also arise from using psychological scales, especially regarding privacy and appropriate representation. Additionally, overfitting to specific traits in the selected scales may limit the model’s generalizability. Addressing these risks requires diversifying data sources and incorporating robust evaluation methods, including human assessments.
@@ Acknowledgements
This work originates from InCharacter introduced by Xintao Wang. The role-playing agents are constructed based on Chat-Haruhi-Suzumiya proposed by Cheng Li. We owe thanks to the early contributors.
Appendix A Psychological Scales
@@
Appendix B Character Selection
@@
Appendix C Evaluation Prompt
We employed various metrics for evaluation. Among them, win-rate and dimensional scoring were directly assessed using a large language model (LLM). The prompts used for these evaluations are listed in Table 4

### v3

@@ Abstract
Role-playing agents (RPA) have been a popular application area for large language models (LLMs), attracting significant interest from both industry and academia. While existing RPAs well portray the characters’ knowledge and tones, they face challenges in capturing their minds, especially for small role-playing language models (RPLMs). In this paper, we propose to enhance RPLMs via personality-indicative data. Specifically, we leverage questions from psychological scales and distill advanced RPAs to generate dialogues that grasp the minds of characters. Experimental results validate that RPLMs trained with our dataset exhibit advanced role-playing capabilities for both general and personality-related evaluations. Code and data are available at https://github.com/alienet1109/RolePersonality.
@@ 1 Introduction
While existing RPLMs well replicate the knowledge and tone of the intended characters, they struggle to capture their minds, in tasks such as personality assessment (Wang et al., 2024a) and decision simulation (Xu et al., 2024). This is partly because existing role-playing datasets focus on character knowledge and tones (Wang et al., 2024b; Shao et al., 2023). However, capturing characters’ minds is crucial for developing authentic RPAs.
In this paper, we propose to develop RPLMs via personality-indicative data. Specifically, we collect these data through questions from psychological scales. These scale questions are designed to quickly capture broad aspects of personality traits in individuals. Hence, we leverage advanced RPAs for the distillation of knowledge from them. Then, we apply these data to develop RPLMs that better capture the minds of the intended characters.
Specifically, we construct a dataset RolePersonality based on questions from 14 different psychological scales, including both single-round and multi-round data, inspired by InCharacter (Wang et al., 2024a). Following the dataset generation, we apply a filtering process using human-annotated personality labels for the selected characters.
@@
Figure 1: The framework of building and utilizing RolePersonality. First, we obtain RolePersonality by distilling advanced RPAs using scale questions. Then, we train RPLMs on RolePersonality to enhance their ability to capture characters’ minds.
@@ 2.1 Role-Playing Language Models
This method utilizes LLMs for generating conversations or human annotations to build and augment datasets. The topics come from the corresponding literature (Shao et al., 2023), general task instructions (Wang et al., 2024b), and special scenarios such as personality tests (Wang et al., 2024a).
@@ 2.2 Construction of Role-Playing Agents
This approach fine-tunes a base model using existing or custom role-playing datasets. Shao et al. (2023); Yu et al. (2024) enhance foundation models with improved role-playing abilities using datasets featuring a variety of characters and scenarios. Zhou et al. (2023); Wang et al. (2024b) tailor LLMs to role-play specific characters.
@@
For more in-depth role-playing of a specific character, many efforts have focused on character-level engineering (Zhou et al., 2023; Wang et al., 2024b). They collect and process character-related data from the corresponding sources, including collecting profile from Wikipedia (Shao et al., 2023). Typically, they add long-term memory to retrieve knowledge about the character based on similarity with the user’s query (Li et al., 2023).
@@ 3.1 Dataset Construction
To simulate the deep thoughts underlying the characters, we generate persona-indicative data by utilizing psychological scale questions, inspired by InCharacter Wang et al. (2024a). In practice, we construct an RPA by pairing the RPLM with descriptions and the memory base of the target character (Li et al., 2023). RPAs are then engaged with open-ended questions derived from established psychological scales. These questions are designed to elicit the character’s mindset and behaviors in various scenarios. The questions were adapted from well-known psychological scales such as the Big Five Inventory (BFI) and 16Personalities. For more details, refer to Sec. B. We start by rewriting psychological scale questions and implementing a selection process to refine the data.
@@
After generating the responses, we implement a filtering process based on the results of the personality assessment. This involves assessing the generated responses to ensure that they align with the intended character’s personality traits. The ground truth comes from human annotations. Responses inconsistent with the character’s personality were discarded to maintain the quality and consistency of the dataset.
@@
Scale Selection
Our scales are sourced from psychological scales (Bem, 1981; Barrick and Mount, 1991), utilizing the questions rewritten by InCharacter (Wang et al., 2024a). However, not all scales are closely related to character personalities. We carefully selected a subset of these scales that best reflect character personality traits, forming the subset Part. The entire set of selected scales constitutes the subset Full.
Question Selection
Not all questions are suitable for all characters. A question that violates the character’s background may induce hallucinations. As a result, we judge whether the question is suitable for the character by llm. Then we exclude questions that do not fit the character’s background.
Multi-Turn Dialogue
We incorporate multi-turn dialogues to maintain conversation consistency and enhance the model’s contextual understanding. We select questions from different dimensions within the same scale. These multi-turn data form subset Multi. The subset consisting of only single-turn data is classified as subset Single.
3.2 Dataset Statistics
4 Experiment
4.1 Settings
Fine-tuning
Baseline
To compare the effectiveness of different datasets, we fine-tune the model with the same settings in three subsets of RolePersonality (Full+Single, Part+Single, Part+Multi), the dataset introduced by CharacterLLM(Shao et al., 2023) and RoleBench(Wang et al., 2024b). For each dataset, we select approximately 20,000 samples for fine-tuning, keeping the data size about the same. These models, along with the original mistral-7B and gpt-3.5-turbo-0301, are subsequently evaluated to assess their performance.
Evaluation Protocols
@@ 4.2 Personality Fidelity (PF)
We use LLMs to judge the character’s personality based on the model’s responses to personality scale questions and compare the judgments with the ground truth, which was determined by human annotators. We selected eight test characters from the dataset proposed by InCharacter (Wang et al., 2024a). All data related to these test characters are excluded from the training set to ensure unbiased evaluation. This metric provides a comprehensive assessment of the model’s ability to accurately reflect the holistic personality traits of a character.
@@ 5 Conclusion
Our approach to developing Role-Playing Language Models (RPLMs) presents several risks. First, reliance on LLM-generated datasets may perpetuate inherent biases and inaccuracies, leading to unintended behaviors. Second, the lack of compliance mechanisms in interview data can result in inconsistencies, undermining authenticity. Third, the absence of human evaluation means subtle nuances in character portrayal may be missed by automated metrics. Ethical concerns also arise from using psychological scales, especially regarding privacy and appropriate representation. Additionally, overfitting to specific traits in the selected scales may limit the generalizability of models trained on different role-playing datasets. Addressing these risks requires diversifying data sources and incorporating robust evaluation methods, including human assessments.
@@ Acknowledgements
This work was supported by the Chinese NSF Major Research Plan(No.92270121). This work originates from InCharacter introduced by Xintao Wang. The role-playing agents are constructed based on Chat-Haruhi-Suzumiya proposed by Cheng Li. We owe thanks to the early contributors. The computations in this research were performed using the CFff platform of Fudan University.
Appendix A Abbreviation Definitions
This section provides the definitions for abbreviations used throughout the paper.
Definition
Subsets
Ful+Sin Subset using all the single-round questions.
Metrics
Acc Accuracy.
Prec Precision.
Rec Recall.
PF Personality Fidelity.
MR Motivation Recognition.
Table 4: Abbreviation and its corresponding definition.
Appendix B Psychological Scales
@@
16P. BFI BSRI CABIN DTDD ECR-R EIS
60 44 60 164 12 36 33
Emp. EPQ-R GSE ICB LMS LOT-R WLEIS
10 100 10 8 9 10 16
Table 5: The question numbers of each scale. 16P refers to 16Personality and Emp. refers to Empathy.
Appendix C Character Selection
@@
We selected characters from a wide variety of sources, covering a broad spectrum of personality types to ensure a well-distributed representation. We demonstrate the 16Personality of characters in table 6
INTJ INTP INFJ INFP
5 1 4 3
ISTJ ISTP ISFJ ISFP
4 2 1 3
ENTJ ENTP ENFJ ENFP
2 5 1 4
ESTJ ESTP ESFJ ESFP
4 2 1 4
Table 6: The personality of selected characters.
Appendix D Multi-turn Dialogue Consistency
To evaluate the ability of models trained on different role-playing datasets in multi-turn dialogue, we assess the personality consistency of its responses across multiple conversational rounds. Specifically, we conduct five rounds of interactions, with the model providing responses to one scale-based question per round. Consistency in evaluations across these rounds reflects the model’s robustness in maintaining personality traits throughout the dialogue. We calculate the scores for each personality dimension in all five rounds, compute the standard deviation for each dimension, and then average these standard deviations across all dimensions. The overall performance is quantified by this final average, with lower scores indicating greater consistency in multi-turn dialogue. The result is demonstrated in table 7, the model fine-tuned with RolePersonality exhibits higher personality concitency during multi-turn dialogues.
Dataset 16 Perosnality BFI
- 13.697 0.4020
RolePersonality 12.796 0.3365
RoleBench 15.060 0.4572
CharacterLLM 14.428 0.4415
Table 7: The average standard deviation of personality assessment during multi-turn dialogues.
Appendix E Evaluation Prompt
We employed various metrics for evaluation. Among them, win-rate and dimensional scoring were directly assessed using a large language model (LLM). The prompts used for these evaluations are listed in Table 8.